OpenStudy (anonymous):

what is the conjugate of a square root a-1?

OpenStudy (misty1212):

could it be \(\sqrt{a}-1\)?

OpenStudy (misty1212):

not really clear could be \(-\sqrt{a}-1\)

OpenStudy (anonymous):

Thank ya ma'am

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (xapproachesinfinity):

hehe do you mean the conjugate of \[\sqrt{a}-1\] it is \[\sqrt{a}+1\]

OpenStudy (triciaal):